# Quorum sensing and absorbing phase transitions in colloidal active matter

### v1

@@ I Introduction
When in a group, microorganisms, plants, animals and humans regulate their dynamics in response to interactions with their neighbors. When walking in a subway station, we adjust our pace depending on the density of the crowd that surrounds us, and organisms as elementary as E. Coli regulate their swimming motion in response to chemicals released by surrounding bacteria. This fundamental feedback mechanism is known as quorum sensing [1] and results in large-scale patterns, wave propagation and biofilm formation, see e.g. [2, 3, 4, 5, 6, 7, 8]. By contrast, the available synthetic active particles are so rudimental that the mesoscopic structure they form hardly ever feedback on their inner motorization [9, 10, 11, 12, 13, 14]. Until now the experimental consequences of multiscale couplings between structure and activity have been only explored using computer-assisted controls of small collections of active units [15, 16, 17, 18].
@@
In this article, we study the phase behavior of large collections of active colloidal rods. First, we establish experimentally that colloidal rods powered by Quincke motors [19] feature quorum sensing capabilities: they can switch their activity on and off depending on their local density. The rods freely roll on solid surfaces in dilute regions, but stand up and arrest their motion in crowded environments. We then experimentally explore the consequences of synthetic quorum sensing by varying independently the activity and packing fraction of our Quincke rods. We show that they first undergo a phase transition where an active isotropic liquid coexists with an arrested phase of standing rods. Further increasing their packing fraction, despite the absence of any change in the energy injection process, all the rods arrest their self-propulsion and form a macroscopic amorphous solid. To elucidate our experimental findings we combine analytical theory and numerical simulations. We show that quorum sensing and steric interactions compete to stabilize an atypical phase separation where flat interfaces supporting a net pressure drop are in mechanical equilibrium. On the one hand, our primitive form of quorum sensing always induces an absorbing phase transition where all the particles freeze their dynamics. On the other hand, steric interactions induce mechanical stresses that can arrest the freezing process and stabilize active liquids.
@@
Altogether our results reveal a novel mechanism for phase coexistence in active matter that could apply to any synthetic or living system where spatial structures emerge from quorum sensing interactions.
@@ II Experiments: engineering quorum-sensing interactions with Quincke rods
[…]
We detail the experimental methods in Appendix A. In short, to investigate the collective dynamics of Quincke rods, we synthesize $3.7\pm 1.5\,\rm\mu m$ long photoresist cylinders of diameter $0.55\pm 0.17\,\rm\mu m$ (aspect ratio: 6.7) [23]. We then turn these inanimate plastic bodies into self-propelled particles using the Quincke motorization principle [24, 12], and observe their collective dynamics over a range of local number density $\rho(\mathbf{r})$ in a V-shape microfluidic channel, see Fig. 1a. Our first control parameter $\rho_{0}$ is the number density of colloids averaged in the observation window, see Fig. 1a. Our second control parameter is the magnitude of the DC voltage which controls the rod activity ($V\in[10\,\rm V,300\,\rm V]$).
@@ II.1 Electrorotation of non-interacting rods
In our experiments, the rods lie on a horizontal solid substrate whose normal direction is aligned with the $\mathbf{E}$ field. As a result, the gravitational and electrophoretic torques acting on the rods destabilize the vertical configurations. We thus never observe static vertical rods in isolation, neither in colloidal rod experiments, nor using a hundred times bigger 3D printed particles, see Fig. 1c and Supplementary Movie 3. When dilute, Quincke rods behave similarly to spherical colloidal rollers when sitting on solid surfaces: they are either static and parallel to the surface at low field, or spin and roll along their short axis when $E>E_{\rm Q}$, see Fig. 1c.
@@ II.2 Phase behavior of interacting Quincke rods
Figure 2: Phase behavior of interacting Quincke rods. (a) Bright field picture of a homogeneous liquid phase ($\rho_{0}=0.12\,\rm\mu m^{-2}$, $V=150\,\rm V$). The color indicates the local orientation of the rods measured using the method of Rezakhaniha et al. [25]. The active liquid is isotropic. Scale bar: $12\,\rm\mu m$. (b) Fluorescence image of the same liquid. Unlike in the gas phase, some rods have switched to a standing state: they stop spinning and point in the vertical direction (bright dots). The shaded circles help locating the arrested rods. Scale bar: $6\,\rm\mu m$. We stress that only $25\,\%$ of the rods are labeled. (c) Average fraction of standing rods $\alpha/(\alpha+\beta)$ plotted against the average density $\rho_{0}$. (d) Fluorescence image of arrested clusters coexisting with an active liquid phase ($\rho_{0}=0.21\,\rm\mu m^{-2}$, $V=80\,\rm V$). In the clusters, the vast majority of the rods stand in the direction of the $\mathbf{E}$ field. Scale bar: $12\,\rm\mu m$. (e) Three subsequent pictures showing the capture and release of active rods at the interface between the arrested and active-liquid phases. Motile rods may switch to the standing state when colliding with the clusters. Similarly, arrested rods can resume their self-propulsion and escape the cluster boundary. Scale bar: $5\,\rm\mu m$. (f) The area fraction of the arrested phase $\Phi_{A}$ grows when $\rho_{0}$ increases ($V=150\,\rm V$). Fitting the coexistence points with a straight line allows us to define the two binodals $\rho_{\rm L}$ ($\Phi_{A}=0)$ and $\rho_{\rm A}$ ($\Phi_{A}=1)$, corresponding to the vertical shaded lines. (g) At the onset of solidification all the $\Phi_{A}(\rho_{0}-\rho_{\rm L})$ collapse on a master curve (where the binodal $\rho_{L}$ is defined in (f), yellow vertical line) (h) Phase diagram of interacting Quincke rods. The three macroscopic regions correspond to the homogeneous active liquid, the homogeneous arrested solid and to the coexistence region.
@@
Although the transition rates from the rolling to the standing state ($\alpha$) and from the standing to the rolling state ($\beta$) cannot be measured individually, the fraction of standing rods is given by $\alpha/(\alpha+\beta)$ and provides a measure of their ratio, see Fig. 2c. Despite the small statistics, we find that $\alpha$ is vanishingly small below a threshold density $\bar{\rho}\simeq 0.1\,\rm\mu m^{-2}$, and increases above this threshold. This behavior implies that the spontaneous transitions between the rolling and the arrested states is the result of interactions. It is not a single-rod property. This seemingly mundane observation is at the core of our work: Quincke rods achieve a primitive form of quorum sensing. Their activity depends on the local number of neighboring particles.
Increasing $\rho_{0}$, or lowering the actuation voltage $V$, a second phase behavior emerges. A finite fraction of rods switches to the arrested state, stand up and form dense solid clusters hosting no inner dynamics, Fig. 2d and Supplementary Movie 4. These clusters of static standing particles coexist with a homogeneous active liquid. We observe a dynamical exchange of particles at their interface, see Fig. 2e and Supplementary Movie 5. Increasing $\rho_{0}$ results in an increase of the area fraction of the arrested phase $\Phi_{A}$ until all the rods stop propelling and form a homogeneous arrested phase.
@@ II.3 Quorum-sensing-induced phase separation
Figure 3: Quorum sensing induces phase separation. (a) Density of the active liquid ($\rho_{\rm L}$) and of the arrested phases ($\rho_{\rm A}$) plotted against the average rod density $\rho_{0}$. Horizontal lines correspond to the binodals determined in Fig. 2f. Voltage: $V=150\,\rm V$. (b) Growth of a a single arrested cluster in a homogeneous active liquid. The typical size $R(t)$ is defined as the square root of the cluster area. $V=109\,\rm V$, $\rho_{0}=0.15\,\rm\mu m^{-2}$. Scale bar: $8\,\rm\mu m$. (c) Growth kinetics showing the classical $R(t)\sim(t-t_{0})^{1/3}$ scaling law. (d) Image sequence showing the coarsening patterns upon a quench deep in the coexistence region, from $V=300\,V$ to $V=90\,V$. $\rho_{0}=0.20\,\rm\mu m^{-2}$. Scale bar: $9\,\rm\mu m$. (e) Hysteresis loops. When the motorization voltage varies periodically the fraction of the arrested phase displays a hysteresis loop. $\rho_{0}=0.2\,\rm\mu m^{-2}$. The two hysteresis cycles corresponds to different periods $T$ of the applied voltage. The voltage varies from $80\,\rm V$ to $270\,\rm V$. (Left): $T=60\,\rm s$. (Right): $T=300\,\rm s$. (f) Corresponding time series of $\Phi_{A}(t)$ and of $V(t)$. ($T=300\,s)$
@@ III Theory
To gain more insight into our experimental findings, we model our Quincke rods as 2D active Brownian particles (ABPs) that can switch between two states: a rolling state ($R$), where particles propel at constant speed $v_{0}$, and a standing state ($S$), where they are immobile. As above, we denote by $\alpha$ the rate at which particles switch from the rolling to the standing state, and by $\beta$ the rate at which standing particles resume their rolling motion in random direction. Since the transition to the rolling state is induced only by collisions with rolling particles, we posit that $\beta$ depends only on the local density of rolling particles, which we denote by $\rho_{R}(\mathbf{r})$ (in practice we compute it within a radius set by the rod length $\ell_{r}$). By contrast, the transition from standing to rolling state is induced by any collision. We therefore expect $\alpha$ to depend on the local density of particles $\rho(\mathbf{r})$. In addition, Fig. 2b shows that $\alpha$ vanishes when $\rho<\bar{\rho}$.
We therefore model the transition rates as:
[…]
@@ III.1 A Motility-Induced Absorbing Phase Transition
Figure 4: Quorum-sensing-induced absorbing phase transition. (a) Streamplot of the mean-field dynamics of Eq (3). The dashed gray lines correspond to the dynamics, while solid lines represent the fixed points. For $\rho_{0}<\bar{\rho}$, the active phase is the only stable attractor (yellow line, $\rho_{S}=0$). At intermediate densities $\bar{\rho}<\rho_{0}<\rho_{c}$, the fixed points correspond to mixed configurations (teal line). At large densities, $\rho_{0}>\rho_{c}$, the static phase is the only stable attractor (black line). (b-d) Snapshots from a simulation of ABPs with density-dependent transition rates (Eqs. (1) and (2)). The golden rectangles represent the rolling rods, while black circles correspond to standing rods. When an arrested cluster nucleates, it progressively absorbs the whole system as the particle current ${\bf J}\propto-\nabla_{{\bf r}}\rho_{R}$. In this panel, the length of the rods corresponds to the interaction radius $\ell_{r}=5$. Numerical details: see Appendix B and SM.
@@
First, we note that homogeneous arrested phases with $\rho_{R}=0$ are linearly unstable when $\alpha_{0}<\beta_{0}$ (See SM). This is incompatible with our experimental observations, we therefore focus solely on the regime $\alpha_{0}>\beta_{0}$. In Fig. 4, we illustrate the dynamics of Eq. (3), and the stable homogeneous phases it predicts. In agreement with our experiments, at small density ($\rho_{0}<\bar{\rho}$) the stable attractors define an active phase where all the particles roll and $\rho_{S}=0$ (yellow line). Similarly, at large densities ($\rho_{0}>\rho_{c}\equiv\alpha_{0}\bar{\rho}/(\alpha_{0}-\beta_{0})$), we predict a fully-static phase where $\rho_{R}=0$ (black line). However, in the intermediate density range ($\bar{\rho}<\rho<\rho_{c}$) the mean field dynamics lead to a stable homogeneous mixture of standing and rolling rods, which we never observe in our experiments.

### v2

# Synthetic quorum sensing and absorbing phase transitions in colloidal active matter

@@ I Introduction
When in a group, microorganisms, plants, animals and humans regulate their dynamics in response to interactions with their neighbors. When walking in a subway station, we adjust our pace depending on the density of the crowd that surrounds us, and organisms as elementary as E. Coli regulate their gene expression and swimming motion in response to chemicals released by surrounding bacteria. This fundamental feedback mechanism is known as quorum sensing  [1]. As in all sensing process, quorum sensing relies on changes in the inner dynamics of active bodies to regulate their dynamics in response to external cues. Quorum sensing has spectacular consequences on the macroscopic organisation of active matter. In bacteria colonies, it results in large-scale patterns, wave propagation and biofilm formation, all determined by the interplay between bacterial activity and structural organization, see e.g. [2, 3, 4, 5, 6, 7, 8]. By contrast, the available synthetic active particles are so rudimentary that they cannot sense their environment, the mesoscopic structure they form hardly ever feedback on their inner motorization and their positions and orientations react passively to forces and torques [9, 10, 11, 12, 13, 14].
While theory and simulation have shown that the interplay between structure and activity leads to robust self-organization pathways [15, 16, 17, 18, 19, 20, 21, 22], their experimental investigations have been limited to computer-assisted control of small collections of active units [23, 24, 25, 26]. Developing scalable active systems capable of sensing their environment and regulating their motorization in response to structural changes remains an open challenge.
@@
In this article, we study the phase behavior of large collections of active colloidal rods. First, we experimentally establish that colloidal rods powered by Quincke motors [27] feature a primitive form of quorum sensing: they actively respond to density variations by switching their internal motor on and off. In dilute regions, the rods roll freely on solid surfaces, whereas in crowded environments, they stand upright and arrest their motion. We then experimentally explore the consequences of synthetic quorum sensing by varying independently the activity and packing fraction of our Quincke rods. We show that they first undergo a phase transition where an active isotropic liquid coexists with an arrested phase of standing rods. Further increasing their packing fraction, despite the absence of any change in the energy injection process, all the rods arrest their self-propulsion and form a macroscopic amorphous solid. To elucidate our experimental findings we combine analytical theory and numerical simulations. We show that quorum sensing and steric interactions compete to stabilize an atypical phase separation where flat interfaces supporting a net pressure drop are in mechanical equilibrium. On the one hand, our primitive form of quorum sensing always induces an absorbing phase transition where all the particles freeze their dynamics. On the other hand, unlike e.g. in transitions between turbulent states  [28, 29, 30, 31], steric interactions induce mechanical stresses that can arrest the absorption process and allow the active and passive phases to coexist.
@@
Altogether our results reveal a novel mechanism for phase coexistence in active matter that could apply to any synthetic or living system where spatial structures emerge from sensing interactions.
Figure 1: Motorizing colloidal rods. (a) (Left) Sketch of our microfluidic setup. We first tilt the device to let the rods sediment and accumulate at the tip of the V-shaped channel. We then place the device in the horizontal direction and perform our observations. (Right) Bright Field image of the SU8 rods at density $\rho_{0}=0.16\,\rm\mu m^{-2}$, and voltage $V=150\,\rm V$. Scale bar: $5\,\rm\mu m$. (b) Quincke instabilities. (i) Electrohydrodynamic instability of a dielectic sphere immersed in a conducting fluid. When applying a DC field, the static configuration is unstable when $E>E_{Q}$, and the sphere spins around a random axis normal to $\mathbf{E}$ [32]. (ii) When the particle is anisotropic, it points in the direction of $\mathbf{E}$ when $E$ is small. The rod is static. At large $E$, the particle aligns in a random direction normal to $\mathbf{E}$ and spins around its long axis. At intermediate fields both states are stable and coexist [33, 34] (iii) In the vicinity of a solid surface, rods that are dense or charged enough are driven toward the surface. Their behavior mirrors that of of Quincke spheres. When $E<E_{Q}$ the rod is static and aligned with the solid surface, above $E_{Q}$ it rolls along the solid wall in the direction of its short axis (c) Superimposed images of SU-8 colloidal rods (length: $3.7\,\rm\mu m$ diameter: $0.55\,\rm\mu m$) sedimented on a solid electrode for $E/E_{Q}=0,\,2$. The $\mathbf{E}$ field points in the direction normal to the images. The color indicates time (Time step $0.1\,\rm s$). The polydispersity in the rod length results in a distribution Quincke thresholds. We define $E_{Q}$ as the field where the first rod starts rolling. Scale bar: $5\,\rm\mu m$.
@@ II Experiments: engineering quorum-sensing interactions with Quincke rods
We detail the experimental methods in Appendix A. In short, to investigate the collective dynamics of Quincke rods, we synthesize $3.7\pm 1.5\,\rm\mu m$ long photoresist cylinders of diameter $0.55\pm 0.17\,\rm\mu m$ (aspect ratio: 6.7) [35]. We then turn these inanimate plastic bodies into self-propelled particles using the Quincke motorization principle [36, 12], and observe their collective dynamics over a range of local number density $\rho(\mathbf{r})$ in a V-shape microfluidic channel, see Fig. 1a. Our first control parameter $\rho_{0}$ is the number density of colloids averaged in the observation window, see Fig. 1a. The smallest $\rho_{0}$ corresponds to well separated rods all lying on the bottom electrode. At the highest $\rho_{0}$ values the rods accumulate over two layers and form a quasi two-dimensional material. Our second control parameter is the magnitude of the DC voltage which controls the rod activity ($V\in[10\,\rm V,300\,\rm V]$).
@@ II.1 Electrorotation of non-interacting rods
In our experiments, the rods lie on a horizontal solid substrate whose normal direction is aligned with the $\mathbf{E}$ field. As a result, the gravitational and electrophoretic torques acting on the rods destabilize the vertical configurations. We thus never observe static vertical rods in isolation, neither in colloidal rod experiments, nor using a hundred times bigger 3D printed particles, see Fig. 1c and Supplementary Movie 3. When dilute, Quincke rods behave similarly to spherical colloidal rollers. When sitting on solid surfaces: they are either static and parallel to the surface at low field, or spin and roll along their short axis when $E>E_{\rm Q}$, see Fig. 1c. We stress that their direction of motion is not slaved to any external field. Isolated Quincke colloids propel autonomously along random directions.
@@ II.2 Phase behavior of interacting Quincke rods
Figure 2: Phase behavior of interacting Quincke rods. (a) Bright field picture of a homogeneous liquid phase ($\rho_{0}=0.12\,\rm\mu m^{-2}$, $V=150\,\rm V$). The color indicates the local orientation of the rods measured using the method of Rezakhaniha et al. [37]. The active liquid is isotropic. (b) Fluorescence image of the same liquid. Unlike in the gas phase, some rods have switched to a standing state: they stop spinning and point in the vertical direction (bright dots). The shaded circles help locating the arrested rods. We stress that only $25\,\%$ of the rods are fluorescent, see Appendix A. (c) Average fraction of standing rods $\alpha/(\alpha+\beta)$ plotted against the average density $\rho_{0}$. (d) Fluorescence image of arrested clusters coexisting with an active liquid phase ($\rho_{0}=0.21\,\rm\mu m^{-2}$, $V=80\,\rm V$). In the clusters, the vast majority of the rods stand in the direction of the $\mathbf{E}$ field. Note that only 25% of the rods are fluorescent, see Appendix A. (e) Three subsequent pictures showing the capture and release of active rods at the interface between the arrested and active-liquid phases. Motile rods may switch to the standing state when colliding with the clusters. Similarly, arrested rods can resume their self-propulsion and escape the cluster boundary. (f) The area fraction of the arrested phase $\Phi_{A}$ grows when $\rho_{0}$ increases ($V=150\,\rm V$). Fitting the coexistence points with a straight line allows us to define the two binodals $\rho_{\rm L}$ ($\Phi_{A}=0)$ and $\rho_{\rm A}$ ($\Phi_{A}=1)$, corresponding to the vertical shaded lines. (g) At the onset of solidification all the $\Phi_{A}(\rho_{0}-\rho_{\rm L})$ collapse on a master curve (where the binodal $\rho_{L}$ is defined in (f), yellow vertical line) (h) Phase diagram of interacting Quincke rods. The three macroscopic regions correspond to the homogeneous active liquid, the homogeneous arrested solid and to the coexistence region.
@@
Although the transition rates from the rolling to the standing state ($\alpha$) and from the standing to the rolling state ($\beta$) cannot be measured individually, the fraction of standing rods is given by $\alpha/(\alpha+\beta)$ and provides a measure of their ratio, see Fig. 2c. Despite the small statistics, we find that $\alpha$ is vanishingly small below a threshold density $\bar{\rho}\simeq 0.1\,\rm\mu m^{-2}$, and increases above this threshold. This behavior implies that the spontaneous transitions between the rolling and the arrested states is not a single-rod property. Instead, it results from the physical couplings between the Quincke rods that are mediated by contact, electrostatic and hydrodynamic interactions. Despite the lack of gene regulation and chemical signaling, the rod activity depend on the local number of neighboring particles. This seemingly mundane observation is at the core of our work: Quincke rods achieve a primitive form of quorum sensing.
Increasing $\rho_{0}$, or lowering the actuation voltage $V$, a second phase behavior emerges from these quorum sensing interactions. A finite fraction of rods switches to the arrested state, stand up and form dense solid clusters hosting no inner dynamics, Fig. 2d and Supplementary Movie 4. These clusters of static standing particles coexist with a homogeneous active liquid. We observe a dynamical exchange of particles at their interface, see Fig. 2e and Supplementary Movie 5. Increasing $\rho_{0}$ results in an increase of the area fraction of the arrested phase $\Phi_{A}$ until all the rods stop propelling and form a homogeneous arrested phase.
@@ II.3 Quorum-sensing-induced phase separation
Figure 3: Quorum sensing induces phase separation. (a) Density of the active liquid ($\rho_{\rm L}$) and of the arrested phases ($\rho_{\rm A}$) plotted against the average rod density $\rho_{0}$. Horizontal lines correspond to the binodals determined in Fig. 2f. Voltage: $V=150\,\rm V$. (b) Growth of a a single arrested cluster in a homogeneous active liquid. The typical size $R(t)$ is defined as the square root of the cluster area. $V=109\,\rm V$, $\rho_{0}=0.15\,\rm\mu m^{-2}$. Scale bar: $5\,\rm\mu m$. (c) Growth kinetics showing the classical $R(t)\sim(t-t_{0})^{1/3}$ scaling law. (d) Image sequence showing the coarsening patterns upon a quench deep in the coexistence region, from V=300VV=300\,V to V=90VV=90\,V. $\rho_{0}=0.20\,\rm\mu m^{-2}$. Scale bar: $5\,\rm\mu m$. (e) Hysteresis loops. When the motorization voltage varies periodically the fraction of the arrested phase displays a hysteresis loop. $\rho_{0}=0.2\,\rm\mu m^{-2}$. The two hysteresis cycles corresponds to different periods $T$ of the applied voltage. The voltage varies from $80\,\rm V$ to $270\,\rm V$. (Left): $T=60\,\rm s$. (Right): $T=300\,\rm s$. (f) Corresponding time series of $\Phi_{A}(t)$ and of $V(t)$. (T=300s)T=300\,s)
@@ III Theory
Building a mathematical model of Quincke rods that accurately captures contact, electrostatic, and hydrodynamic interactions remains beyond the reach of current theories and simulations. To gain insight, we therefore construct a robust minimal model that isolates the essential ingredients needed to explain—and generalize—our experimental findings.
We model our Quincke rods as 2D active Brownian particles (ABPs) that can switch between two states: a rolling state ($R$), where particles propel at constant speed $v_{0}$, and a standing state ($S$), where they are immobile. As above, we denote by $\alpha$ the rate at which particles switch from the rolling to the standing state, and by $\beta$ the rate at which standing particles resume their rolling motion in random direction. Since the transition to the rolling state is induced only by collisions with rolling particles, we posit that $\beta$ depends only on the local density of rolling particles, which we denote by $\rho_{R}(\mathbf{r})$ (in practice we compute it within a radius set by the rod length $\ell_{r}$). By contrast, the transition from standing to rolling state is induced by any collision. We therefore expect $\alpha$ to depend on the local density of particles $\rho(\mathbf{r})$. In addition, Fig. 2b shows that $\alpha$ vanishes when $\rho<\bar{\rho}$.
Informed by our observations and measurements, we hence model the transition rates as:
$R\xrightarrow{\alpha}S\;,$ $\quad\alpha({\rho})\>\>\>=\alpha_{0}({\rho}-\bar{\rho})\Theta({\rho}-\bar{\rho})$ (1)
$S\xrightarrow{\beta}R\;,$ $\quad\beta({\rho}_{R})=\beta_{0}{\rho}_{R}\;,$ (2)
@@ III.1 A Motility-Induced Absorbing Phase Transition
Figure 4: Quorum-sensing-induced absorbing phase transition. (a) Streamplot of the mean-field dynamics of Eq (3). The dashed gray lines correspond to the dynamics, while solid lines represent the fixed points. For $\rho_{0}<\bar{\rho}$, the active phase is the only stable attractor (yellow line, $\rho_{S}=0$). At intermediate densities $\bar{\rho}<\rho_{0}<\rho_{c}$, the fixed points correspond to mixed configurations (teal line). At large densities, $\rho_{0}>\rho_{c}$, the static phase is the only stable attractor (black line). (b-d) Snapshots from a simulation of ABPs with density-dependent transition rates (Eqs. (1) and (2)). To facilitate comparison between experiments and simulations, we plot the rolling rods as golden rectangles and the standing rods as black circles. We show in SM that simulations of rods interacting via actual anisotropic quorum-sensing interactions lead to identical results. When an arrested cluster nucleates, it progressively absorbs the whole system as the particle current ${\bf J}\propto-\nabla_{{\bf r}}\rho_{R}$. In this panel, the length of the rods corresponds to the interaction radius $\ell_{r}=5$. Numerical details: see Appendix B and SM.
@@
First, we note that homogeneous arrested phases with $\rho_{R}=0$ are linearly unstable when $\alpha_{0}<\beta_{0}$ (See SM [43]). This is incompatible with our experimental observations, we therefore focus solely on the regime $\alpha_{0}>\beta_{0}$. In Fig. 4, we illustrate the dynamics of Eq. (3), and the stable homogeneous phases it predicts. In agreement with our experiments, at small density ($\rho_{0}<\bar{\rho}$) the stable attractors define an active phase where all the particles roll and $\rho_{S}=0$ (yellow line). Similarly, at large densities ($\rho_{0}>\rho_{c}\equiv\alpha_{0}\bar{\rho}/(\alpha_{0}-\beta_{0})$), we predict a fully-static phase where $\rho_{R}=0$ (black line). However, in the intermediate density range ($\bar{\rho}<\rho<\rho_{c}$) the mean field dynamics lead to a stable homogeneous mixture of standing and rolling rods, which we never observe in our experiments.